OpenStudy (nayo98):

Please help! I dont even know how to go about this question. How many moles of ozone does the atmosphere hold? For a sense of scale, compare the amount to the total carbondioxide emitted annually from fossil fuel burning, cement production and deforestation (7900 Tg/C). Given with the facts the amount of ozone is equivalent to a 3mm band around earth, standard temperature and pressure is at 22.4 dm^3 with the radius of earth 6378km.

OpenStudy (mww):

Use the gas law \[pV = nRT\] now since they specify standard temperature and pressure has been applied, you can assume then that V = 22.4 dm^3 or 22.4 L/mol (i.e. one mole of any gas at STP occupies 22.4L) So you want to find n the number of moles of ozone, which requires knowledge of V. The volume can be found be considering the volume of a sphere and doing a volume subtraction: (this is assuming the atmosphere is within the radius, if it is beyond, then change r values accordingly so that you add to the radius of the earth.  \[V = \frac{ 4 }{ 3 } \pi r_{outer}^3 - \frac{ 4 }{ 3 } \pi r_{inner} ^3 = \frac{ 4 }{ 3 }\pi(r_{outer}^3 - r_{inner}^3)\] = 1.53 x 10^12 m^3 = 1.53 x 10^15 dm^3 or L. (Note 1 m^3 = 10dm x 10dm x 10dm = 1000 dm^3 or L) The neat thing is now you have a volume measurement in litres. you can find the number of moles by considering that each mole consists of 22.4 L.

OpenStudy (nayo98):

Omg! Thank you so much. I now know what to do